OpenStudy (anonymous):

Simplify: 3x√2•√14x?

OpenStudy (lgbasallote):

\[3x\sqrt 2 \times \sqrt{14x} \implies 3x\sqrt{2 \times 14x} \implies 3x \sqrt{28x} \implies 3x \sqrt{4 \times 7x}\]

OpenStudy (lgbasallote):

\[\implies 3x \times 2 \sqrt{7x}\]

OpenStudy (lgbasallote):

sorry for the lack of space